OpenStudy (anonymous):

you invest an initial $4,500 in an account that has an annual interest rate of4.5% compounded daily. How much money will you have in the account after 10 years? Round your answer to the nearest whole number.

OpenStudy (anonymous):

"P" stands for principle. "R" stands for rate or interest rate "M*" stands for months "T" stands for time so it says, "you invest an initial $4,500 in an account," which means money you are spending now or how much you start with. P=4500[1+.045/(compound daily)]^(m**10)

OpenStudy (anonymous):

the formula is: P2=P1*(1-r/m)^(m*t)

OpenStudy (anonymous):

it is saying compound daily, which means 365 days. it is said monthly, it m=12 months quarterly, m=4 get it? so, P=4500*(1+.045/365)^(365*10) <-- you can just put this in the cal and get your final answer.

OpenStudy (anonymous):

I dont think Im getting the right answer?

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

180

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

Really, thanks a lot you really helped me.

OpenStudy (anonymous):

i wrote the formula sign wrong. it's A=P*(1+r/m)^(m*t)